# Optimising quantum tomography via shadow inversion

### v2

@@ Introduction
Quantum state tomography with classical shadows Huang et al. (2020), initially motivated by the seminal work of Aaronson Aaronson (2017), represents a significant advancement in the realm of quantum information processing, particularly in the efficient characterisation of quantum systems. This technique, developed as a solution to the prohibitive resource requirements of full quantum state tomography has since seen a plethora of modifications Hu et al. (2023); Elben et al. (2020); Huang et al. (2021); Neven et al. (2021); Crawford et al. (2021); Chen et al. (2021); Hadfield et al. (2022); Zhao et al. (2021); Seif et al. (2023); Wan et al. (2023); Levy et al. (2024); Bu et al. (2024); Coopmans et al. (2023); Bertoni et al. (2023); Koh and Grewal (2022); McGinley and Fava (2023); Kunjummen et al. (2023); Iosue et al. (2024); Nakaji et al. (2023); Feldman et al. (2022); Wyderka et al. (2023), which all seek to provide a viable alternative for obtaining meaningful information about quantum states with substantially reduced computational and experimental overhead Sack et al. (2022); Anshu et al. (2021); Huang et al. (2023); Nguyen et al. (2022); Zhang et al. (2021); Struchalin et al. (2021); Stilck França et al. (2024); Ippoliti et al. (2023); Morris et al. (2022); Glos et al. (2022); Stricker et al. (2022); Vermersch et al. (2023). Naturally, the main focus has been the ’quantum’ element of the procedure as this is where most practical efforts are stymied, either in gate or sample complexity. Other than machine learning Huang et al. (2023, 2022) and adaptive techniques García-Pérez et al. (2021); Glos et al. (2022), little thought has been put towards explicit optimisation of the classical postprocessing with it often being treated as a mostly fixed step.
@@
for $T_{k\ell}=U_{k}^{\dagger}\Pi_{\ell}U_{k}$. The inverse operation $\mathcal{M}^{-1}$ may be easily computed to recover the input state $\rho$ from the output $\mathcal{M}^{-1}$. Classical shadows’ great insight was to show that this inversion procedure remains effective even when only a tiny fraction of Clifford measurements are used in the above sum. This rapid convergence to a known configuration means the linear map of (2) may be readily inverted to recover the original state $\rho$. As shown in Huang et al. (2020), the inverse map of (2) for global Clifford measurements on $N$ qubits seems to easily appear as the inverse depolarising channel: $\mathcal{M}_{N}^{-1}(X)=(2^{N}+1)X-\mbox{Tr}\left[{X}\right]\mathbbm{1}_{2^{N}}$; in the case of tensor products of random single-qubit Pauli measurements, the corresponding inverted quantum channel reduces to $\mathcal{M}^{-1}_{\mathcal{P}}=\bigotimes_{i=1}^{N}\mathcal{M}_{1}^{-1}$.
@@
$\expectationvalue{\hat{\eta}}=\sum_{k=1}^{n}p_{k}\hat{\eta}_{k}=\sum_{k=1}^{n}\mbox{Tr}\left[{\bar{\rho}E_{k}}\right]\hat{\eta}_{k}\overset{!}{=}\bar{\rho},~{}~{}~{}\forall\bar{\rho}\in\mathcal{V}_{D}.$ (3)
This is the requirement on the estimator $\hat{\eta}$ such that the statistics of the shadow mimics, in expectation, that of the actual unknown state. Since the POVM forms an over-complete basis in $\mathcal{V}_{D}$, any solution to Eq. (3) is not unique. This multiplicity results in a set of free parameters, which becomes an extra resource that may be used in the post-processing component of an estimation to reduce the variance.
To achieve this, we now show a constructive method to recover families of equivalent classical shadows in terms of the POVM elements and their free parameters. We rely on an equivalent problem to (3), that is, to impose the equivalence between the expectation value of an observable $A$ and the corresponding classical estimator $a=\mbox{Tr}\left[{A\hat{\eta}}\right]$:
$\expectationvalue{\,a\,}=\mbox{Tr}\left[{A\expectationvalue{\hat{\eta}}}\right]=\sum_{k=1}^{n}p_{k}\mbox{Tr}\left[{A\,\hat{\eta}_{k}}\right]\overset{!}{=}\mbox{Tr}\left[{\bar{\rho}A}\right].$ (4)
It is instructive to express the operators as vectors in $\mathcal{V}_{D}$, i.e., for an operator $X\in\mathcal{V}_{D}$, we shall use the ket notation $|{X}\rangle\!\rangle$. We also identify an orthonormal basis $\beta_{D}=\{|{B_{j}}\rangle\!\rangle\}_{j=1}^{D}$ in $\mathcal{V}_{D}$. The expectation value from (4) can be expressed as
$\expectationvalue{\,a\,}=\sum_{k=1}^{n}\langle\!\langle{A^{\dagger}}|{\eta_{k}}\rangle\!\rangle\langle\!\langle{E_{k}^{\dagger}}|{\rho}\rangle\!\rangle\overset{!}{=}\langle\!\langle{A^{\dagger}}|{\bar{\rho}}\rangle\!\rangle$ (5)
where $\langle\!\langle{X}|{Y}\rangle\!\rangle=\mbox{Tr}\left[{X^{\dagger}Y}\right]$. To guarantee the estimators are unbiased, the requirement reduces to a resolution of the identity in the subspace $\mathcal{V}_{D}$
@@
$\sum_{k=1}^{n}\langle\!\langle{B_{s}}|{\hat{\eta}_{k}}\rangle\!\rangle\langle\!\langle{E_{k}}|{B_{t}}\rangle\!\rangle=\sum_{k=1}^{n}L_{sk}R_{tk}^{*}\overset{!}{=}\delta_{st},$ (7)
with $L_{sk}=\langle\!\langle{B_{s}}|{\hat{\eta}_{k}}\rangle\!\rangle$ and $R_{tk}=\langle\!\langle{B_{t}}|{E_{k}}\rangle\!\rangle$. This implies the matrix equation $LR^{\dagger}=\mathbbm{1}_{D}$; thus, $L$ is a generalised inverse of $R^{\dagger}$. Since $R$ is a $D\times n$ matrix having $|{E_{k}}\rangle\!\rangle$ as columns, it is a rank-$D$ matrix (recall there are only $D$ independent $E_{k}$ matrices). We can write its singular value decomposition (SVD), i.e., $R=U(\Sigma|\hat{0})W^{\dagger}$ with $\Sigma=\mathrm{diag}[\sigma_{1},\dots,\sigma_{D}]$ ($\sigma_{s}>0$) and $\hat{0}$ being the $D\times(n-D)$ zero-matrix. Having this, a simple exercise shows $L=W(\Sigma^{-1}|H)U^{\dagger}$ with $H$ being an arbitrary $D\times(n-D)$ matrix. Accordingly, we can split solution into two parts $L=L^{(p)}+L^{(h)}$, with $L^{(p)}=W(\Sigma^{-1}|\hat{0})U^{\dagger}$ (particular) and $L^{(h)}=W(\hat{0}|H)U^{\dagger}$ (homogeneous). This results in a family of equivalent estimators
@@
parameterised by a homogeneous solution $\hat{\eta}_{k}^{(h)}$. As we shall see, this multiplicity of possible inversion schemes represents an additional resource for error optimisation in post-processing. We turn to observable properties which may be expressed in terms of estimator coefficients $a_{k}=\mbox{Tr}\left[{A\hat{\eta}_{k}}\right]=a_{k}^{(p)}+a_{k}^{(h)}$, with $a_{k}^{(p/h)}=\mbox{Tr}\left[{A\hat{\eta}_{k}^{(p/h)}}\right]$ where the homogeneous coefficients do not influence expectation values or any other linear combination of classical shadows
@@
By substituting $p_{k}=\mbox{Tr}\left[{\rho E_{k}}\right]$ and $\expectationvalue{\,a\,}=\mbox{Tr}\left[{AE_{k}}\right]$ into the previous equation we find
@@
Figure 2: (a) Optimal free parameter for projectors on the Bloch sphere. The contour plot shows the minimal shadow norm (13) of the single-qubit projector characterised by the Bloch sphere angles $\theta,\,\phi$ (in units of $\pi$). Optimisation results in bounds lower or equal than $3/2$, obtained via the Clifford inversion map (15) found in Huang et al. (2020) (b) Optimal free parameter for plane projectors. The optimal value of the shadow norm of the single-qubit projector in the equatorial plane in terms of a single angle parameter $\phi$ (in units of $\pi$). This enhanced norm is always lower or equal to the standard shadow norm |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}\leq 5/4$ (see main text). In the case of $\phi=k\frac{\pi}{2},\,k\in\mathbb{Z}$, which corresponds to the projectors on the Pauli $X$ and $Y$ basis. The bound corresponds to the combinations which present optimal variance |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}\equiv 1$.
Now, to effectively compare the different choices for estimators, we refer to the efficiency of the estimation process, that is the expected sample size of copies required to estimate a property of the state up to a fixed accuracy. It is customary in classical shadow tomography to refer to variance as figure of merit Huang et al. (2020). In particular, we consider the state-independent upper bound on variance, defined as the shadow norm, which only depends on the decomposition of the particular observable:
$\mbox{Var}\left[{a}\right]\leq\max_{\sigma:\text{state}}\mbox{Tr}\left[{\sigma\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}}\right]=\|A\|^{2}_{\text{shadow}}.$ (12)
This quantity is equivalent to the maximal eigenvalue of the variance operator defined as $\hat{O}_{A}\coloneq\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}$; we will use both definitions interchangeably. This is akin to the original upper bound on shadow norm and thence sampling complexity found in Huang et al. (2020), where the expected number of samples scales with $M=O(\varepsilon^{-2}\log{\delta^{-1}}\|A\|^{2}_{\text{shadow}})$ ). Nevertheless, this can be further optimised over the free parameters in the homogeneous part of the estimator. By optimising over these coefficients, it is possible to identify the configuration which guarantees the minimal sample size for the estimation of the particular observable $A$:
@@
The existence of a minimum is guaranteed through the positive semidefiniteness of the variance operator. We will provide examples to illustrate that different choices of free parameters may result in a sampling complexity with exponential separation compared to that of the naive estimator construction, which shows, in turn, that optimisation in (13) yields exceptional gains for relatively minuscule cost.
@@
The optimisation in Eq. (12) appears efficient, given the total number of free parameters scales linearly with the number of subsystems $N$. We see that even minimal deviation from the minimum in local shadow norms $\|A_{j}\|^{2}_{\text{opt}}$ may accumulate exponentially fast with $N$. This is illustrated in Figure 1(b) for the case of the canonical inversion map of Eq. (15). We now present examples of such exponential improvement for qubits for particular sets of local observables, specifically projectors, whereby by adjusting the available free parameters, we are able to reduce the expected bound on sample size based on assumptions for the observable itself.
@@
We first consider the case of estimation of fidelity with product states. The target observables are thus tensor products of a projector on a single qubit $\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}$ |, parameterised on the Bloch sphere by two angles $(\theta_{j},\,\phi_{j})$. The measurement scheme is described by the POVM of projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{2}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y,\,z$. As reference, we consider the bound $\|\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{shadow}}\equiv(3/2)^{N}$ obtained using the estimators
$\hat{\eta}^{\pm}_{\zeta}=3\left(3E^{\pm}_{\zeta}-\mbox{Tr}\left[{E^{\pm}_{\zeta}}\right]\mathbbm{1}_{3}\right).$ (15)
@@
found in Huang et al. (2020). Now we show how to improve this bound by our techniques. The POVMs $E^{\pm}_{\zeta}$ span the entire HS space. We have $n=6$ and $D=4$, thus the equation (11) gives two free optimisation parameters for each qubit See supplementary material at [URL from publisher] for information concerning estimator explicit parameter dependence . In Fig.2(a), we present the minimal upper bound as a function of angles $(\theta,\phi)$, which is contained within the interval $1.15\leq\|\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{sh}}\leq 1.5$. Since the overall shadow norm (14) is of the product form, there are entire regions in the $(\theta,\phi)$-space for which deviations from from the standard $3/2$ leads to the exponential improvement in $N$.
@@
We now present a configuration in which the optimisation of the inversion map not only improves the standard bound, but we are actually able to achieve efficiency, i.e. $\|A\|^{2}_{\text{opt}}\equiv 1$. We consider the POVM constituted by the projectors of the X and Y Paulis, which spans the ”equator” on the Bloch sphere. The target observables are tensor products of a projector on a single qubit, parameterised by a single angle: $A=\outerproduct{\psi(\phi)}{\psi(\phi)}$ | with $|{\psi(\phi)}\rangle=(|{0}\rangle+e^{i\phi}|{1}\rangle)\sqrt{2}$. As before, we follow the ansatz given in Huang et al. (2020) to get the particular solution for the inversion map and estimators
From here we obtain the reference bound for the shadow norm of $\|\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{sh}}\equiv(5/4)^{N}$. We reference to Supplementary Information See supplementary material at [URL from publisher] for information concerning shadow norm expression for local Clifford measurements for more precise remarks on how we obtain this result.
Now we turn to our optimisation technique. We have the POVM with $n=4$ elements spanning a subspace of dimension $D=3$, therefore the generic estimators obtained from our procedure depend on a single free parameter $p$. The shadow norm of single-qubit projector depends on this free parameter and on the angle parameter $\phi$:
@@
In general, the dependency on $\phi$ remains explicit, changing the expected bound based on the projector to be estimated. In Fig.2(b) we present the optimal value of the shadow norm for each parameter $\phi\in[0,\pi]$. We note that for each value of $\phi$ the optimised norm is |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{opt}}\leq 5/4$. Moreover, it can be seen that for specific angles, the bound is exactly $\|E_{\zeta}^{\pm}\|^{2}_{\text{opt}}\equiv 1$., i.e., for the eigenprojectors of $X$ and $Y$ Paulis, corresponding to the POVM elements themselves. This implies efficient estimation: given the product form of the overall shadow norm (14), this scaling represents not only an exponential improvement compared to $(5/4)^{N}$, as for other projectors, but also an expected constant scaling of the sample size with the number of qubits. This is particularly relevant considering that X and Y are non-commuting observables, as it opens the possibility of the simultaneous estimation of up to $2^{N}$ projections simultaneously for complementary observables. Previously this result has been found only via global measurement strategies relying on Clifford unitaries consisting of many joint operations, acting on all $N$ qubits at the same time; we have shown instead that by introducing the initial assumption over the space of observables, this could be achieved using only local single-qubit measurements.

### v3

@@ Introduction
Quantum state tomography with classical shadows Huang et al. (2020), initially motivated by the seminal work of Aaronson Aaronson (2017), represents a significant advancement in the realm of quantum information processing, particularly in the efficient characterisation of quantum systems. This technique, developed as a solution to the prohibitive resource requirements of full quantum state tomography has since seen a plethora of modifications Hu et al. (2023); Elben et al. (2020); Huang et al. (2021a); Neven et al. (2021); Crawford et al. (2021); Chen et al. (2021); Hadfield et al. (2022); Zhao et al. (2021); Seif et al. (2023); Wan et al. (2023); Levy et al. (2024); Bu et al. (2024); Coopmans et al. (2023); Bertoni et al. (2023); Koh and Grewal (2022); McGinley and Fava (2023); Kunjummen et al. (2023); Iosue et al. (2024); Nakaji et al. (2023); Feldman et al. (2022); Wyderka et al. (2023), which all seek to provide a viable alternative for obtaining meaningful information about quantum states with substantially reduced computational and experimental overhead Sack et al. (2022); Anshu et al. (2021); Huang et al. (2023); Nguyen et al. (2022); Zhang et al. (2021); Struchalin et al. (2021); Stilck França et al. (2024); Ippoliti et al. (2023); Morris et al. (2022); Glos et al. (2022); Stricker et al. (2022); Vermersch et al. (2023). Naturally, the main focus has been the ’quantum’ element of the procedure as this is where most practical efforts are stymied, either in gate or sample complexity. Other than machine learning Huang et al. (2023, 2022) and adaptive techniques García-Pérez et al. (2021); Glos et al. (2022), little thought has been put towards explicit optimisation of the classical postprocessing with it often being treated as a mostly fixed step.
@@
for $T_{k\ell}=U_{k}^{\dagger}\Pi_{\ell}U_{k}$. The inverse operation $\mathcal{M}^{-1}$ may be easily computed to recover the input state $\rho$ from the output $\mathcal{M}^{-1}$. Classical shadows’ great insight was to show that this inversion procedure remains effective even when only a tiny fraction of Clifford measurements are used in the above sum. This rapid convergence to a known configuration means the linear map of Eq. (2) may be readily inverted to recover the original state $\rho$. As shown in Huang et al. (2020), the inverse map of Eq. (2) for global Clifford measurements on $N$ qubits seems to easily appear as the inverse depolarising channel: $\mathcal{M}_{N}^{-1}(X)=(2^{N}+1)X-\mbox{Tr}\left[{X}\right]\mathbbm{1}_{2^{N}}$; in the case of tensor products of random single-qubit Pauli measurements, the corresponding inverted quantum channel reduces to $\mathcal{M}^{-1}_{\mathcal{P}}=\bigotimes_{i=1}^{N}\mathcal{M}_{1}^{-1}$.
@@
$\expectationvalue{\hat{\eta}}=\sum_{k=1}^{n}p_{k}\hat{\eta}_{k}=\sum_{k=1}^{n}\mbox{Tr}\left[{\bar{\rho}E_{k}}\right]\hat{\eta}_{k}=\bar{\rho},~{}~{}~{}\forall\bar{\rho}\in\mathcal{V}_{D}.$ (3)
This is the requirement on the estimator $\hat{\eta}$ such that the statistics of the shadow mimics, in expectation, that of the actual unknown state. With Eq. (3), the task of finding an inversion map can be made equivalent to identifying a dual frame to the POVM in space of operators Innocenti et al. (2023). Since the POVM forms an over-complete basis in $\mathcal{V}_{D}$, the solution to Eq. (3) is not unique. This multiplicity results in a set of free parameters, which becomes an extra resource that may be used in the post-processing component of an estimation to reduce the variance.
To achieve this, we now show a constructive method to recover families of equivalent classical shadows in terms of the POVM elements and their free parameters. We rely on an equivalent problem to Eq. (3), that is, to impose the equivalence between the expectation value of an observable $A$ and the corresponding classical estimator $a=\mbox{Tr}\left[{A\hat{\eta}}\right]$:
$\expectationvalue{\,a\,}=\mbox{Tr}\left[{A\expectationvalue{\hat{\eta}}}\right]=\sum_{k=1}^{n}p_{k}\mbox{Tr}\left[{A\,\hat{\eta}_{k}}\right]=\mbox{Tr}\left[{\bar{\rho}A}\right].$ (4)
It is instructive to express the operators as vectors in $\mathcal{V}_{D}$, i.e., for an operator $X\in\mathcal{V}_{D}$, we shall use the ket notation $|{X}\rangle\!\rangle$. We also identify an orthonormal basis $\beta_{D}=\{|{B_{j}}\rangle\!\rangle\}_{j=1}^{D}$ in $\mathcal{V}_{D}$.
The expectation value from Eq. (4) can be expressed as
$\expectationvalue{\,a\,}=\sum_{k=1}^{n}\langle\!\langle{A^{\dagger}}|{\eta_{k}}\rangle\!\rangle\langle\!\langle{E_{k}^{\dagger}}|{\rho}\rangle\!\rangle=\langle\!\langle{A^{\dagger}}|{\bar{\rho}}\rangle\!\rangle$ (5)
where $\langle\!\langle{X}|{Y}\rangle\!\rangle=\mbox{Tr}\left[{X^{\dagger}Y}\right]$ is the inner product in $\mathcal{V}_{D}$.
To guarantee the estimators are unbiased, the requirement reduces to a resolution of the identity in the subspace $\mathcal{V}_{D}$
@@
$\sum_{k=1}^{n}\langle\!\langle{B_{s}}|{\hat{\eta}_{k}}\rangle\!\rangle\langle\!\langle{E_{k}}|{B_{t}}\rangle\!\rangle=\sum_{k=1}^{n}L_{sk}R_{tk}^{*}=\delta_{st},$ (7)
with $L_{sk}=\langle\!\langle{B_{s}}|{\hat{\eta}_{k}}\rangle\!\rangle$ and $R_{tk}=\langle\!\langle{B_{t}}|{E_{k}}\rangle\!\rangle$. This implies the matrix equation $LR^{\dagger}=\mathbbm{1}_{D}$; thus, $L$ is a generalised inverse of $R^{\dagger}$. Since $R$ is a $D\times n$ matrix having $|{E_{k}}\rangle\!\rangle$ as columns, it is a rank-$D$ matrix (recall there are only $D$ independent $E_{k}$ matrices). We can write its singular value decomposition (SVD), i.e., $R=U(\Sigma|\hat{0})W^{\dagger}$ with $\Sigma=\mathrm{diag}[\sigma_{1},\dots,\sigma_{D}]$ ($\sigma_{s}>0$) and $\hat{0}$ being the $D\times(n-D)$ zero-matrix. Having this, a simple exercise shows $L=U(\Sigma^{-1}|H)W^{\dagger}$ with $H$ being an arbitrary $D\times(n-D)$ matrix. Accordingly, we can split solution into two parts $L=L^{(p)}+L^{(h)}$, with $L^{(p)}=U(\Sigma^{-1}|\hat{0})W^{\dagger}$ (particular) and $L^{(h)}=U(\hat{0}|H)W^{\dagger}$ (homogeneous). This results in a family of equivalent estimators
@@
parameterised by a homogeneous solution $\hat{\eta}_{k}^{(h)}$. As we shall see, this multiplicity of possible inversion schemes represents an additional resource for error optimisation in post-processing. We turn to observable properties which may be expressed in terms of estimator coefficients $a_{k}=\mbox{Tr}\left[{A\hat{\eta}_{k}}\right]=a_{k}^{(p)}+a_{k}^{(h)}$, with $a_{k}^{(p/h)}=\mbox{Tr}\left[{A\hat{\eta}_{k}^{(p/h)}}\right]$. The homogeneous coefficients, from the definition of $\hat{\eta}_{k}^{(h)}$ in Eq. (8), do not influence expectation values or any other linear combination of classical shadows
@@
By substituting $p_{k}=\mbox{Tr}\left[{\rho\,E_{k}}\right]$ and $\expectationvalue{\,a\,}=\mbox{Tr}\left[{A\,\rho}\right]$ into the previous equation we find
@@
Figure 2: (a) Optimal free parameter for projectors on the Bloch sphere. The contour plot shows the minimal shadow norm Eq. (13) of the single-qubit projector characterised by the Bloch sphere angles $\theta,\,\phi$ (in units of $\pi$). Optimisation results in bounds lower or equal than $3/2$, obtained via the Clifford inversion map Eq. (15) found in Huang et al. (2020) (b) Optimal free parameter for plane projectors. The optimal value of the shadow norm of the single-qubit projector in the equatorial plane in terms of a single angle parameter $\phi$ (in units of $\pi$). This enhanced norm takes values in the range $41\leq\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}\leq 5/$ 4 (see main text). In the case of $\phi=k\frac{\pi}{2},\,k\in\mathbb{Z}$, which corresponds to the projectors on the Pauli $X$ and $Y$ basis, the bound achieves optimal value |$\|\outerproduct{\psi}{\psi}(\phi)\|^{2}_{\text{sh}}\equiv 1$.
Now, to effectively compare the different choices for estimators, we refer to the efficiency of the estimation process, that is the expected sample size of copies required to estimate a property of the state up to a fixed accuracy. It is customary in classical shadow tomography to refer to variance as figure of merit Huang et al. (2020). In particular, we consider the state-independent upper bound on variance, called shadow norm $\|A\|^{2}_{\text{sh}}$, defined as:
$\begin{split}\mbox{Var}\left[{a}\right]&=\mathbb{E}[a^{2}]-(\mathbb{E}[a])^{2}\leq\mathbb{E}[a^{2}]=\sum_{k=1}^{n}p_{k}|a_{k}|^{2}\\ &\leq\max_{\sigma:\text{state}}\mbox{Tr}\left[{\sigma\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}}\right]=\|A\|^{2}_{\text{sh}}.\end{split}$ (12)
Considering the upper bound to study sample complexity is reasonable since, for bounded observables, the quantity $(\mathbb{E}[a])^{2}$ is also bounded, therefore the relevant contribution is brought by the first term $\mathbb{E}[a^{2}]$.
This quantity is equivalent to the maximal eigenvalue of the variance operator defined as $\hat{O}_{A}\coloneq\sum_{k=1}^{n}|a_{k}^{(p)}+a_{k}^{(h)}|^{2}E_{k}$; we will use both definitions interchangeably. This is akin to the original upper bound on shadow norm and thence sampling complexity found in Huang et al. (2020), where the expected number of samples scales with $M=O(\varepsilon^{-2}\log{\delta^{-1}}\|A\|^{2}_{\text{sh}})$ ). Nevertheless, this can be further optimised over the free parameters in the homogeneous part of the estimator. By optimising over these coefficients, it is possible to identify the configuration which guarantees the minimal sample size for the estimation of the particular observable $A$:
@@
The existence of a minimum is guaranteed through the positive semidefiniteness of the variance operator. We will provide examples to illustrate that different choices of free parameters may result in a sampling complexity with exponential separation compared to that of the naive estimator construction, which shows, in turn, that optimisation in Eq. (13) yields exceptional gains for relatively minuscule cost.
@@
Since we are considering local measurements on independent subsystems, each term $\|A_{j}\|^{2}_{\text{opt}}$ can be optimised locally, implying the optimisation in Eq. (14) is efficient overall. We see that even minimal deviation from the minimum in local shadow norms $\|A_{j}\|^{2}_{\text{opt}}$ may accumulate exponentially fast with $N$. This is illustrated in Figure 1(b) for the case of the canonical inversion map of Eq. (15). We now present examples of such exponential improvement for qubits for particular sets of local observables, specifically projectors, whereby by adjusting the available free parameters, we are able to reduce the expected bound on sample size based on assumptions for the observable itself.
@@
We first consider the case of estimation of fidelity with product states. The target observables are thus tensor products of a projector on a single qubit $\bigotimes_{j=1}^{N}\outerproduct{\psi_{j}}{\psi_{j}}$ |, parameterised on the Bloch sphere by two angles $(\theta_{j},\,\phi_{j})$. The measurement scheme is described by the POVM of normalised projectors on the Pauli basis $E^{\pm}_{\zeta}=\frac{1}{6}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y,\,z$. Using the estimators
$\hat{\eta}^{\pm}_{\zeta}=3\left(3E^{\pm}_{\zeta}-\mbox{Tr}\left[{E^{\pm}_{\zeta}}\right]\mathbbm{1}_{2}\right).$ (15)
found in Huang et al. (2020), the shadow norm in Eq. (12) simply evaluates for the generic single-qubit observable $A=a\cdot\mathbbm{1}+x\cdot\sigma_{x}+y\cdot\sigma_{y}+z\cdot\sigma_{z}$ to
$\|A\|^{2}_{\text{sh}}=a^{2}+3(x^{2}+y^{2}+z^{2})+2|a|\sqrt{x^{2}+y^{2}+z^{2}}.$ (16)
Considering the particular case of projectors on the Bloch sphere, with the parameterisation
$\begin{split}a\to\frac{1}{2},\quad\quad x\to\frac{1}{2}\sin\theta\cos\phi,\\ y\to\frac{1}{2}\sin\theta\sin\phi,\quad\quad z\to\frac{1}{2}\cos\theta,\end{split}$ (17)
and Eq. (16) trivially reduces to |$\|\outerproduct{\psi}{\psi}\|^{2}_{\text{sh}}\equiv\frac{3}{2}$. Now we show how to improve this reference bound by our techniques. The POVMs $E^{\pm}_{\zeta}$ span the entire HS space and we may determine easily a homogeneous combination of the effects as in Eq. (11) given $\sum_{\zeta}h_{\zeta}^{+}E_{\zeta}^{+}+h_{\zeta}^{-}E_{\zeta}^{-}=0$. Exploiting the symmetry of the system, they reduce to three parameters since $h_{\zeta}^{+}=h_{\zeta}^{-}$:
$\{h_{\zeta}\}_{\zeta=x,y,z}=\{p_{1},p_{2},-p_{1}-p_{2}\}$ (18)
As expected, we find $n-D=2$ parameters $p_{1},p_{2}$.
@@
Optimising over these parameters, we present the minimal upper bound as a function of angles $(\theta,\phi)$ in Fig.2(a). The optimised results are contained within the interval $1.15\leq\|\outerproduct{\psi_{j}}{\psi_{j}}\|^{2}_{\text{sh}}\leq 1.5$. Since the overall shadow norm Eq. (14) is of the product form, there are entire regions in the $(\theta,\phi)$-space for which deviations from from the standard $3/2$ leads to the exponential improvement in $N$.
Estimation with constrained POVMs
In the previous section we have studied examples of overcomplete POVMs in the entire HS space, i.e. with $D=d^{2}$. Here we will pursue a relevant example of a POVM spanning only a subspace $\mathcal{V}_{D}$ of lower dimension $D<d^{2}$, in which $E_{k}$ forms an over-complete basis (within $\mathcal{V}_{D}$). Note that the standard inversion procedure Huang et al. (2020); Nguyen et al. (2022) is defined on the full HS space and does not directly apply in this instance.
Consider for instance the POVM constituted by the projectors of the X and Y Paulis $E^{\pm}_{\zeta}=\frac{1}{4}(\mathbbm{1}\pm\sigma_{\zeta})$, where $\zeta=x,\,y$, which spans the ”equator” on the Bloch sphere. With a straightforward application of our procedure Eq. (8), we obtain the following particular solution
The homogeneous term is added by solving $\sum_{\zeta}h_{\zeta}^{+}E_{\zeta}^{+}+h_{\zeta}^{-}E_{\zeta}^{-}=0$, where $\{h_{\zeta}\}_{\zeta=x,y}=\{p,-p\}$ and $p$ is a free parameter . For the generic planar observable (within span of $\{E_{\zeta}^{\pm}\}$)
$A=\matrixquantity(a&x+iy\\ x-iy&a),$ (20)
we find the shadow norm
$\|A\|^{2}_{\text{sh}}(p)$ $=(a-x)^{2}+2(x^{2}+y^{2})+B(p)+\sqrt{C(p)},$ (21)
$B(p)$ $=\frac{1}{8}(2(p_{r}^{2}+p_{i}^{2})a^{2}+(p_{r}(8+p_{r})+p_{i}^{2})x^{2}$
$+(p_{r}(p_{r}-8)+p_{i}^{2})y^{2}),$
$C(p)$ $=x^{2}(a(8+p_{r}(6+p_{r})+p_{i}^{2})-2(4+p_{r})x)^{2}$
$+(a(8+(p_{r}-6)p_{r}+p_{i}^{2})+2(p_{r}-4)x)^{2}y^{2},$
[…]
We can now proceed to two examples of estimation of product observables via local measurements achieving the same scaling of resources as global strategies Huang et al. (2020).
Optimal estimation of planar Paulis
When considering as target observable the planar Pauli,
$a\to 0,\quad x\to\cos\varphi,\quad y\to\sin\varphi,$ (22)
the optimised shadow norm from Eq. (21) results in $\|A\|^{2}_{\text{opt}}=2$. An immediate result is that the expected number of resources for product of planar Paulis scales as $2^{N}$, which matches the scaling obtained from global Clifford measurements Huang et al. (2020), $\|A\|^{2}_{\text{sh}}=\mbox{Tr}\left[{A^{2}}\right]=2^{N}$. Surprisingly, we are able to match the optimal scaling Huang et al. (2021b) while still relying only on local single-qubit measurements.
Complete versus overcomplete POVMs
In the context of the previous example, it is interesting to compare the performance of complete versus overcomplete POVM. For this purpose, consider the example of a triangle POVM on the plane, defined as the normalised projectors $E_{k}=\frac{2}{3}\outerproduct{\psi_{k}}{\psi_{k}}$ | of states
$|{\psi_{k}}\rangle=\dfrac{|{0}\rangle+e^{i\frac{k}{3}2\pi}|{1}\rangle}{\sqrt{2}}\quad k=0,1,2$ (23)
Simple inversion of this complete basis brings us to estimators
$\hat{\eta}_{k}=3E_{k}-\frac{3}{4}\mbox{Tr}\left[{E_{k}}\right]\mathbbm{1}_{2}.$ (24)
Since the POVM is complete, there is no free parameter dependence. From here, we can compute the shadow norm as defined in Eq.(12). For a planar Pauli as target observable, as in Eq. (22), the shadow norm is $\|A\|^{2}_{\text{sh}}=3,\,\forall\varphi$, which leads to an expected overhead of $3^{N}$ on the scaling of resources. From the previous example, the same observables can be optimally estimated with an overhead of $2^{N}$. We have therefore found an example in which an overcomplete POVM performs better than a complete measurement scheme in the same subspace.
@@
We now present a configuration in which the optimisation of the inversion map not only improves the sample complexity, but is actually able to achieve efficient estimation, i.e. $\|A\|^{2}_{\text{opt}}=1$.
To show this, we choose as target observables tensor products of projectors that act on a qubit, $A=\bigotimes_{j=1}^{N}\outerproduct{\psi(\phi_{j})}{\psi(\phi_{j})}$ | where each single-qubit state is parameterised as: $|{\psi(\phi_{j})}\rangle=(|{0}\rangle+e^{i\phi_{j}}|{1}\rangle)\sqrt{2}$.
Using the change of coordinates
$a\to\frac{1}{2},\quad x\to\frac{1}{2}\cos\phi,\quad y\to\frac{1}{2}\sin\phi,$ (25)
we express the generic shadow norm from Eq. (21) in terms of the parameter $\phi$:
@@
This is particularly relevant considering that X and Y are non-commuting observables, as it opens the possibility of the simultaneous estimation of up to $4^{N}$ projections simultaneously for complementary observables. Previously this result has been found in Huang et al. (2020) only via global measurement strategies relying on $N-$qubit Clifford unitaries. We have shown instead that this is achieved using exclusively local measurements.